# Target Speaker Extraction through Comparing Noisy Positive and Negative Audio Enrollments

### v2

@@ Abstract
Target speaker extraction focuses on isolating a specific speaker’s voice from an audio mixture containing multiple speakers. To provide information about the target speaker’s identity, prior works have utilized clean audio samples as conditioning inputs. However, such clean audio examples are not always readily available. For instance, obtaining a clean recording of a stranger’s voice at a cocktail party without leaving the noisy environment is generally infeasible. Limited prior research has explored extracting the target speaker’s characteristics from noisy enrollments, which may contain overlapping speech from interfering speakers. In this work, we explore a novel enrollment strategy that encodes target speaker information from the noisy enrollment by comparing segments where the target speaker is talking (Positive Enrollments) with segments where the target speaker is silent (Negative Enrollments). Experiments show the effectiveness of our model architecture, which achieves over 2.1 dB higher SI-SNRi compared to prior works in extracting the monaural speech from the mixture of two speakers. Additionally, the proposed two-stage training strategy accelerates convergence, reducing the number of optimization steps required to reach 3 dB SNR by 60%. Overall, our method achieves state-of-the-art performance in the monaural target speaker extraction conditioned on noisy enrollments.
@@ 1 Introduction
In the target speaker extraction task, the model is required to extract the target speaker’s voice from a mixture of multiple speakers and ambient noise. To specify the characteristics of the target speaker, prior works have explored using conditional information of the target speaker in multiple modalities, including visual pan2022selectivelistening ; muse , contextual ma2024clapsep ; typingtolistern , or acoustic modality speakerbeam ; spex ; spex+ ; zhang2024multilevelspeakerrepresentationtarget . Though prior works using conditional information in the acoustic modality have achieved significant good extraction performance, most of these works only considered using clean audio examples of the target speaker he2024hierarchicalspeaker ; meng2024binauralselective ; zhao2024continuoustargetspeech ; pham2024wannahearvoiceadaptive . This strong assumption prevents these models from performing well when only noisy audio examples are available. For example, consider a cocktail party, where the user meets a stranger who has not provided an audio sample before. To extract the target stranger speaker’s voice from the noisy environment to assist in clear conversation, the user will have to ask the speaker to step outside to record the clean audio enrollment. Enrolling target speakers in this way is often impractical in real-world applications.
@@ 3.3 Model Architecture
The Positive and Negative Enrollments are first encoded by a pair of TF-GridNet encoders in the Siamese Encoder. Since The goal of both branches in the Siamese Encoder is to encode multiple speakers’ characteristics from the noisy Positive or Negative Enrollments, we share the parameters between the two TF-GridNet encoders of the Siamese Encoder to reduce model parameter size. The Siamese encoder results in two sequences of embeddings $E_{pos}$ and $E_{neg}$ with shapes $[T_{pos},D]$ and $[T_{neg},D]$, respectively.
The Encoder Fusion Module then extracts the target speaker’s voice embedding from the Siamese Encoders’ output $E_{pos}$ and $E_{neg}$. To encode the target speaker’s characteristic that exists throughout the Positive Enrollment but is absent in the Negative Enrollment, we need to perform a pairwise comparison between frame-level embeddings of both enrollments. First, we element-wise add each frame embedding in both sequences with learnable segmentation embeddings, to indicate whether it originates from the Positive or Negative Enrollment. We then adopt a self-attention mechanism, which allows the model to dynamically compute similarity scores between all frame embeddings. These attention scores serve as a soft alignment, enabling each positive frame to attend to other positive frame embeddings to remove the Positive Interferers’ voice characteristic from its encoding, and attend to the negative frame embeddings to remove Negative Interferers’ voice characteristics. After two Full-band Self-attention Modules, the model is able to selectively encode only target-speaker-specific features while suppressing interfering signals from background noise or other speakers. We retain only the embeddings corresponding to the positive enrollment as the extracted target speaker embedding, which reduces the output to a $[T_{\text{pos}},D]$ shaped embedding $E_{\text{fused}}$. This reduction in embedding size ensures the embedding feature is the same size as the pretrained encoder’s output, allowing the model to benefit from the fast convergence brought by training the encoder using knowledge distillation (as shown in Section 4.3). To further reduce the computation load in the extraction branch, we apply non-overlapping average pooling with 40 kernel size along the temporal dimension to $E_{\text{fused}}$, resulting in a $[T_{\text{pos}}//40,D]$ shaped feature $E_{\text{fused-pooled}}$. We discuss the effect of different kernel sizes on the model performance and inference time in Appendix D.
@@ 4.1 Datasets and Baselines
We compare our monaural model’s performance with TCE Chen2024tce , SpeakerBeam speakerbeam , and non-negative matrix factorization (NMF) method NMF . We compare our binaural model’s reverberant target speech extraction performance with the LookOnceToHear model Veluri2024lookonce .
@@
3. SpeakerBeam speakerbeam is a target speech extraction model using clean target speaker enrollment as the extraction condition. We include this model as a baseline to show the influence of having interfering speakers in the audio enrollment on the model performance.
@@ 4.2 Result on Monaural and Binaural Reverberant Target Speech Extraction
Method Audio Type Metric 2 Speakers in Extraction 3 Speakers in Extraction
@@ 2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
NMF NMF Mono SNRi 4.24$\pm$1.60 5.65$\pm$1.85
SI-SNRi -1.65$\pm$3.78 -1.50$\pm$3.57
STOI 0.362$\pm$0.139 0.296$\pm$0.114
DNSMOS 1.56$\pm$0.44 1.55$\pm$0.41
SpeakerBeam speakerbeam Mono SNRi 4.78$\pm$2.87 4.84$\pm$2.83 4.84$\pm$2.85 5.91$\pm$2.52 5.58$\pm$2.48 5.72$\pm$2.48
SI-SNRi 2.49$\pm$12.19 1.75$\pm$11.89 0.51$\pm$11.86 1.11$\pm$8.46 -0.33$\pm$8.70 -1.07$\pm$8.84
STOI 0.617$\pm$0.271 0.595$\pm$0.270 0.564$\pm$0.268 0.486$\pm$0.212 0.451$\pm$0.212 0.440$\pm$0.213
DNSMOS 1.72$\pm$0.57 1.73$\pm$0.59 1.71$\pm$0.61 1.57$\pm$0.51 1.51$\pm$0.40 1.49$\pm$0.52
TCE Chen2024tce Mono SNRi 8.48$\pm$2.34 7.84$\pm$2.56 7.52$\pm$2.68 8.47$\pm$2.37 8.57$\pm$2.34 8.07$\pm$2.40
SI-SNRi 6.67$\pm$3.69 5.57$\pm$4.47 4.77$\pm$5.34 6.63$\pm$3.74 5.13$\pm$4.07 4.10$\pm$4.60
STOI 0.682$\pm$0.120 0.650$\pm$0.140 0.624$\pm$0.160 0.682$\pm$0.120 0.551$\pm$0.143 0.521$\pm$0.151
DNSMOS 1.85$\pm$0.44 1.84$\pm$0.44 1.83$\pm$0.44 1.85$\pm$0.44 1.65$\pm$0.41 1.65$\pm$0.41
Mono SNRi 8.53$\pm$2.30 8.39$\pm$2.37 8.37$\pm$2.50 9.03$\pm$2.23 9.00$\pm$2.31 8.89$\pm$2.38
Ours SI-SNRi 6.70$\pm$3.43 6.45$\pm$3.74 6.31$\pm$3.84 6.25$\pm$3.35 6.06$\pm$3.68 5.76$\pm$3.95
(Film fusion) STOI 0.681$\pm$0.111 0.675$\pm$0.123 0.665$\pm$0.127 0.581$\pm$0.123 0.578$\pm$0.129 0.565$\pm$0.133
DNSMOS 1.92$\pm$0.39 1.90$\pm$0.39 1.90$\pm$0.41 1.69$\pm$0.38 1.68$\pm$0.38 1.67$\pm$0.38
Mono SNRi 10.14$\pm$2.57 9.93$\pm$2.71 9.79$\pm$2.86 10.42$\pm$2.49 10.37$\pm$2.64 10.22$\pm$2.79
Ours SI-SNRi 8.85$\pm$3.67 8.54$\pm$4.01 8.30$\pm$4.47 8.42$\pm$3.62 8.29$\pm$4.06 7.82$\pm$4.62
(Monaural) STOI 0.758$\pm$0.107 0.749$\pm$0.121 0.742$\pm$0.126 0.668$\pm$0.123 0.665$\pm$0.130 0.648$\pm$0.146
DNSMOS 2.14$\pm$0.37 2.13$\pm$0.37 2.02$\pm$0.38 1.93$\pm$0.37 1.92$\pm$0.38 1.90$\pm$0.38
Bi SNRi 9.36$\pm$3.56 8.90$\pm$3.57 9.12$\pm$3.57 9.30$\pm$3.39 9.28$\pm$3.49 9.30$\pm$3.56
LookOnceToHear SI-SNRi 7.75$\pm$5.35 7.24$\pm$5.43 7.53$\pm$5.35 6.58$\pm$5.37 6.49$\pm$5.62 6.51$\pm$5.64
Veluri2024lookonce STOI 0.736$\pm$0.152 0.723$\pm$0.162 0.732$\pm$0.153 0.620$\pm$0.175 0.610$\pm$0.180 0.611$\pm$0.187
DNSMOS 1.79$\pm$0.56 1.77$\pm$0.55 1.76$\pm$0.55 1.59$\pm$0.47 1.60$\pm$0.48 1.58$\pm$0.49
Bi SNRi 9.84$\pm$3.57 9.60$\pm$3.57 9.59$\pm$3.63 9.81$\pm$3.28 9.78$\pm$3.23 9.83$\pm$3.30
Ours SI-SNRi 8.18$\pm$4.81 7.84$\pm$5.20 7.75$\pm$5.17 6.76$\pm$5.27 6.72$\pm$5.08 6.71$\pm$5.50
(Binaural) STOI 0.735$\pm$0.150 0.729$\pm$0.155 0.720$\pm$0.159 0.608$\pm$0.177 0.605$\pm$0.177 0.605$\pm$0.182
DNSMOS 1.82$\pm$0.60 1.80$\pm$0.59 1.80$\pm$0.59 1.59$\pm$0.50 1.60$\pm$0.49 1.59$\pm$0.48
@@
We compare our model with baseline methods under scenarios where different numbers of speakers are present. Table 1 shows the monaural target speaker extraction performance. SpeakerBeam speakerbeam shows significantly worse performance, especially as the number of interfering speaker increase in the enrollment. TCE Chen2024tce achieves better performance than SpeakerBeam speakerbeam baseline since it does not rely on the clean target speaker’s enrollment. However, apart from the STOI metric in one scenario, the TCE model shows worse performance than our method under all different numbers of interfering speakers in the Audio Mixture and in enrollments.
@@
In the multi-channel target speech extraction, prior work Veluri2024lookonce has explored using beamforming to extract the target speaker’s characteristics. In this experiment, we show that additionally including the Negative Enrollment helps improve model performance. As shown in Table 1, in comparison to the LookOnceToHear Veluri2024lookonce baseline, based on two-sample t-tests over the 5000 test samples at the 90% confidence level, our model achieves statistically significant improvements in SNRi and SI-SNRi under all conditions. The worse STOI performance might be caused by the difference in parameter size, as shown in Appendix C. In contrast to the monaural experiment results, the binaural models show lower performance in most of the metrics and scenarios. This performance drop might be caused by the difference in training objective, since the binaural models need to predict the reverberant binaural audio, while the monaural models are trained to extract non-reverberant speech (to achieve fair comparison with the prior works focusing on non-reverberant extraction). The DNSMOS metric is also affected by the reverberation effect in the prediction, thus providing less effective information for comparing model performance. We include it here only for completeness.
@@ 4.3 Ablation Study
Such a performance difference might be caused by the high difficulty in encoding the expected target speaker from the noisy enrollment pair. Since its very challenging for the model encoder to encode the target speaker from the noisy enrollment pair, it is difficult for the model to draw similarity between its extraction ground-truth and the noisy pair of enrollment input, resulting in the end-to-end learning providing little guidance to the encoder on which speaker is expected to be encoded. In contrast, the two-stage training explicitly guides the model to extract the target speaker’s embedding in the first stage, significantly accelerating convergence.

### v3

@@ Abstract
Target speaker extraction focuses on isolating a specific speaker’s voice from an audio mixture containing multiple speakers. To provide information about the target speaker’s identity, prior works have utilized clean audio samples as conditioning inputs. However, such clean audio examples are not always readily available. For instance, obtaining a clean recording of a stranger’s voice at a cocktail party without leaving the noisy environment is generally infeasible. Limited prior research has explored extracting the target speaker’s characteristics from noisy enrollments, which may contain overlapping speech from interfering speakers. In this work, we explore a novel enrollment strategy that encodes target speaker information from the noisy enrollment by comparing segments where the target speaker is talking (Positive Enrollments) with segments where the target speaker is silent (Negative Enrollments). Experiments show the effectiveness of our model architecture, which achieves over 2.1 dB higher SI-SNRi compared to prior works in extracting the monaural speech from the mixture of two speakers. Additionally, the proposed two-stage training strategy accelerates convergence, reducing the number of optimization steps required to reach 3 dB SNR by 60%. Overall, our method achieves state-of-the-art performance in the monaural target speaker extraction conditioned on noisy enrollments. Our implementation is available at https://github.com/xu-shitong/TSE-through-Positive-Negative-Enroll .
@@ 1 Introduction
In the target speaker extraction task, the model is required to extract the target speaker’s voice from a mixture of multiple speakers and ambient noise. To specify the characteristics of the target speaker, prior works have explored using conditional information of the target speaker in multiple modalities, including visual pan2022selectivelistening; muse, contextual ma2024clapsep; typingtolistern, or acoustic modality speakerbeam; spex; spex+; zhang2024multilevelspeakerrepresentationtarget. Though prior works using conditional information in the acoustic modality have achieved significant good extraction performance, most of these works only considered using clean audio examples of the target speaker he2024hierarchicalspeaker; meng2024binauralselective; zhao2024continuoustargetspeech; pham2024wannahearvoiceadaptive. This strong assumption prevents these models from performing well when only noisy audio examples are available. For example, consider a cocktail party, where the user meets a stranger who has not provided an audio sample before. To extract the target stranger speaker’s voice from the noisy environment to assist clear conversation, the user will have to ask the speaker to step outside to record the clean audio enrollment. Enrolling target speakers in this way is often impractical in real-world applications.
@@ 3.3 Model Architecture
The Positive and Negative Enrollments are first encoded by a pair of TF-GridNet encoders in the Siamese Encoder. Since The goal of both branches in the Siamese Encoder is to encode multiple speakers’ characteristics from the noisy Positive or Negative Enrollments, we share the parameters between the two TF-GridNet encoders of the Siamese Encoder to reduce model parameter size. The Siamese encoder results in two sequences of embeddings $E_{pos}$ and $E_{neg}$ with shapes $T_{pos},D]$ and $T_{neg},D]$, respectively. $T_{pos},T_{neg}$ are the numbers of frames in the positive and negative enrollment embeddings, and $D$ is the feature dimension at each frame.
# Input:
#   E_pos: shape [T_pos, D]
#   E_neg: shape [T_neg, D]
# Learnable parameters:
#   Segment Embeddings:
#     S_pos, S_neg: shape [1, D]
#   M: 2 Full-band Self-attention layers
Encoder Fusion Module:
  # Step 1: Elementwise add via broadcasting
  E_pos = E_pos + S_pos # shape [T_pos, D]
  E_neg = E_neg + S_neg # shape [T_neg, D]
  # Step 2: Concatenate along the temporal dimension
  E_concat = [E_pos, E_neg] # shape [T_pos + T_neg, D]
  # Step 3: Apply two Full-band Self-attention layers
  E_concat = M(E_concat) # shape [T_pos + T_neg, D]
  # Step 4: Truncate embeddings
  E_fused = E_concat[:T_pos] # shape [T_pos, D]
Figure 3: Encoder Fusion Module pseudo-code.
The Encoder Fusion Module then extracts the target speaker’s voice embedding from the Siamese Encoders’ output $E_{pos}$ and $E_{neg}$. The pseudo-code for the Encoder Fusion Module is shown in Figure 3. In Step 1, two segment embeddings, $S_{pos}$ and $S_{neg}$ with shape $[1,D]$, are first element-wise added to the input embeddings to allow the model to distinguish which enrollment each embedding originates from. The resulting two embeddings are concatenated along the temporal dimension in Step 2, and passed through two Full-band Self-attention layers in Step 3. In these two layers, the self-attention calculation between the embeddings of different Positive Enrollment frames allows the model to identify the Positive Interferers, who remain silent in some of the Positive Enrollment frames, and exclude these speakers in the extracted embedding. Similarly, attention between the Positive and Negative Enrollments embeddings enables the model to identify and exclude Negative Interferers, (i.e., speakers present only in the negative enrollments), from the output embedding. Finally, in Step 4, we truncate the output to retain only the embeddings corresponding to the positive enrollment as the extracted target speaker embedding, which reduces the output to a $T_{\text{pos}},D]$ shaped embedding $E_{\text{fused}}$. This reduction in embedding size ensures the embedding feature is the same size as the pretrained encoder’s output, allowing the model to benefit from the fast convergence brought by training the encoder using knowledge distillation (as shown in Section 4.3). To further reduce the computation load in the extraction branch, we apply non-overlapping average pooling with 40 kernel size along the temporal dimension to $E_{\text{fused}}$, resulting in a $T_{\text{pos}}//40,D]$ shaped feature $E_{\text{fused-pooled}}$. We discuss the effect of different kernel sizes on the model performance and inference time in Appendix D.
@@
In addition to the monaural model described above, which processes single-channel audio, we propose a binaural variant designed to encode and extract binaural audios. The difference between the monaural and binaural model architectures lies solely in the input. The binaural model takes in 4 channels input (stacked real and imaginary components from both channels) and the monaural model takes in only 2 channels (stacked real and imaginary component). By stacking the two channels’ STFT, the binaural model can capture the inter-channel temporal differences between the two channels, which encode the directional information of the target speaker since the target speaker speaks at 90 degree azimuth angle in the enrollment. Both model architectures do not differ after the first convolution layer.
@@ 4.1 Datasets and Baselines
We compare our monaural model’s performance with TCE Chen2024tce, USEF-TFGridnet zeng2024useftse, and non-negative matrix factorization (NMF) method NMF. We compare our binaural model’s reverberant target speech extraction performance with the LookOnceToHear model Veluri2024lookonce.
@@
3. USEF-TFGridnet zeng2024useftse is a target speech extraction model using clean target speaker enrollment as the extraction condition. We include this model as a baseline to show the influence of having interfering speakers in the audio enrollment on the model performance.
@@ 4.2 Result on Monaural and Binaural Reverberant Target Speech Extraction
Method Audio Type Metric 2 Speakers in Mixture 3 Speakers in Mixture
@@ 2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
[…]
USEF-TFGridnet zeng2024useftse Mono SNRi 3.42 $\pm$ 3.43 3.45 $\pm$ 3.58 3.30 $\pm$ 3.50 4.31 $\pm$ 2.52 4.15 $\pm$ 2.47 4.23 $\pm$ 2.58
SI-SNRi -0.03 $\pm$ 5.97 -0.03 $\pm$ 6.42 -0.14 $\pm$ 6.00 0.29 $\pm$ 3.38 0.11 $\pm$ 3.36 0.09 $\pm$ 3.24
PESQ 1.52 $\pm$ 0.49 1.54 $\pm$ 0.49 1.52 $\pm$ 0.49 1.32 $\pm$ 0.38 1.32 $\pm$ 0.37 1.31 $\pm$ 0.36
STOI 0.43 $\pm$ 0.17 0.43 $\pm$ 0.18 0.43 $\pm$ 0.17 0.36 $\pm$ 0.11 0.35 $\pm$ 0.11 0.36 $\pm$ 0.11
DNSMOS 1.37 $\pm$ 0.45 1.37 $\pm$ 0.45 1.36 $\pm$ 0.45 1.35 $\pm$ 0.41 1.33 $\pm$ 0.41 1.34 $\pm$ 0.42
WER 0.66 $\pm$ 0.33 0.66 $\pm$ 0.36 0.68 $\pm$ 0.32 0.85 $\pm$ 0.26 0.86 $\pm$ 0.21 0.86 $\pm$ 0.20
[…]
PESQ 1.91 $\pm$ 0.34 1.80 $\pm$ 0.41 1.73 $\pm$ 0.46 1.91 $\pm$ 0.34 1.53 $\pm$ 0.39 1.48 $\pm$ 0.40
[…]
WER 0.73 $\pm$ 0.16 0.76 $\pm$ 0.24 0.77 $\pm$ 0.24 0.73 $\pm$ 0.25 0.88 $\pm$ 0.21 0.88 $\pm$ 0.17
[…]
Ours PESQ 1.84 $\pm$ 0.32 1.83 $\pm$ 0.34 1.84 $\pm$ 0.33 1.57 $\pm$ 0.31 1.56 $\pm$ 0.32 1.54 $\pm$ 0.35
[…]
WER 0.54 $\pm$ 0.37 0.54 $\pm$ 0.29 0.54 $\pm$ 0.28 0.73 $\pm$ 0.24 0.72 $\pm$ 0.25 0.73 $\pm$ 0.24
[…]
Ours PESQ 2.07 $\pm$ 0.34 2.06 $\pm$ 0.36 2.05 $\pm$ 0.36 1.79 $\pm$ 0.33 1.78 $\pm$ 0.34 1.75 $\pm$ 0.37
[…]
WER 0.42 $\pm$ 0.35 0.43 $\pm$ 0.28 0.45 $\pm$ 0.28 0.61 $\pm$ 0.28 0.61 $\pm$ 0.28 0.62 $\pm$ 0.28
@@
We compare our model with baseline methods under scenarios where different numbers of speakers are present. Table 1 shows the monaural target speaker extraction performance. USEF-TFGridnet zeng2024useftse shows significantly worse performance, especially as the number of interfering speaker increase in the enrollment. TCE Chen2024tce achieves better performance than USEF-TFGridnet zeng2024useftse baseline since it does not rely on the clean target speaker’s enrollment. However, apart from the STOI metric in one scenario, the TCE model shows worse performance than our method under all different numbers of interfering speakers in the Audio Mixture and in enrollments.
Table 2: Comparison between our method and the baselines’ performance when extracting binaural audio from mixtures of 2-3 speakers conditioned on an enrollment mixture of 2-4 speakers.
Method Audio Type Metric 2 Speakers in Mixture 3 Speakers in Mixture
2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
in Enroll in Enroll in Enroll in Enroll in Enroll in Enroll
[…]
LookOnceToHear PESQ 2.28 $\pm$ 0.47 2.25 $\pm$ 0.52 2.28 $\pm$ 0.45 1.87 $\pm$ 0.48 1.88 $\pm$ 0.44 1.90 $\pm$ 0.47
[…]
WER 0.45 $\pm$ 0.41 0.44 $\pm$ 0.35 0.47 $\pm$ 0.68 0.66 $\pm$ 0.36 0.63 $\pm$ 0.33 0.64 $\pm$ 0.33
[…]
WER 0.45 $\pm$ 0.48 0.45 $\pm$ 0.34 0.44 $\pm$ 0.35 0.63 $\pm$ 0.33 0.64 $\pm$ 0.43 0.64 $\pm$ 0.42
@@
In the multi-channel target speech extraction, prior work Veluri2024lookonce has explored using beamforming to extract the target speaker’s characteristics. In this experiment, we show that additionally including the Negative Enrollment helps improve model performance. As shown in Table 2, in comparison to the LookOnceToHear Veluri2024lookonce baseline, based on two-sample t-tests over the 5000 test samples at the 90% confidence level, our model achieves statistically significant improvements in SNRi and SI-SNRi under all conditions. The worse STOI performance might be caused by the difference in parameter size, as shown in Appendix C. In contrast to the monaural experiment results, the binaural models show lower performance in most of the metrics and scenarios. This performance drop might be caused by the difference in training objective, since the binaural models need to predict the reverberant binaural audio, while the monaural models are trained to extract non-reverberant speech (to achieve fair comparison with the prior works focusing on non-reverberant extraction). The DNSMOS, PESQ, and WER metric are also affected by the reverberation effect in the prediction, thus providing less effective information for comparing model performance. We include them here only for completeness.
@@ 4.3 Ablation Study
Such a performance difference might be caused by the high difficulty in encoding the expected target speaker from the noisy enrollment pair. Since its very challenging for the model to draw similarity between the ground-truth single speaker audio and the noisy pair of enrollments input, training using an end-to-end learning method provides little guidance to the encoder on which speaker is expected to be encoded. In contrast, the two-stage training explicitly guides the model to extract the target speaker’s embedding in the first stage, significantly accelerating convergence.